### v3

# A statistically consistent measure of semantic uncertainty using Language Models

Abstract
To address the challenge of quantifying uncertainty in the outputs generated by language models, we propose a novel measure of semantic uncertainty, semantic spectral entropy, that is statistically consistent under mild assumptions. This measure is implemented through a straightforward algorithm that relies solely on standard, pretrained language models, without requiring access to the internal generation process. Our approach imposes minimal constraints on the choice of language models, making it broadly applicable across different architectures and settings. Through comprehensive simulation studies, we demonstrate that the proposed method yields an accurate and robust estimate of semantic uncertainty, even in the presence of the inherent randomness characteristic of generative language model outputs.
[…]
Appendix
A statistically consistent measure of semantic uncertainty using Language Models
Yi Liu Prorata.ai Bellevue, Washington, USA yi@prorata.ai
1 Introduction
.. The birth of large language models (LLMs) has given rise to the possibility of a wide range of industry applications Touvron et al. (2023); Chowdhery et al. (2023). One of the key applications of generative models that has garnered significant interest is the development of specialized chatbots with domain-specific expertise such as legal and healthcare. These applications illustrate how generative models can improve decision-making and improve the efficiency of professional services in specialized fields.
A significant challenge hindering the widespread industrial deployment of generative models is the lack of consistency in LLM outputs across multiple runs given the same input Amodei et al. (2016); Hendrycks et al. (2021). Without delving into the intricate mathematical architecture of generative language models, one can observe that these models function as stochastic generators: they take an input string and produce an output string. Prior studies Wang et al. (2020, 2023); Song et al. (2024) have demonstrated that the output for a fixed input can vary across runs, with the degree of variation influenced by factors such as temperature, top-k sampling, top-p sampling, and repetition penalties.
In a recent study, Atil et al. (2024) Atil et al. (2024) conducted experiments using six deterministically configured LLMs—each with temperature set to 0 and top-p set to 1—across eight common tasks, with five identical trials per task. The goal was to assess output repeatability under conditions designed to minimize randomness. Surprisingly, none of the models exhibited complete consistency in their outputs across all tasks. For more complex tasks, such as those involving college-level mathematics, the models frequently produced lexically distinct outputs across trials, resulting in zero consistency with respect to exact string matching. However, the authors also observed that many of these outputs were semantically equivalent despite their lexical differences, suggesting that syntactic variation does not necessarily imply divergence in meaning.
Variations in observed responses have been attributed to the use of GPUs during large language model (LLM) inference, where premature rounding and non-deterministic computation can introduce discrepancies Nvidia (2024); Atil et al. (2024). Consequently, it is reasonable to regard LLM outputs as inherently stochastic, with the elimination of this randomness being practically challenging—if not infeasible. This poses a significant concern for the commercial deployment of LLMs: when errors occur (i.e., when an output deviates from the expected result given a specific input), it becomes difficult to disentangle implementation-induced randomness from genuine model deficiencies. This ambiguity complicates efforts to make targeted improvements to the system. In response to this challenge, we propose the need for an empirical measure of uncertainty that evaluates the semantic variability of LLM outputs, rather than relying on surface-level lexical comparisons. Such a measure can be interpreted as a variance estimator that captures the semantic dispersion within a set of generated outputs.
Most prior studies on uncertainty in foundation models for natural language processing (NLP) have focused primarily on the calibration of classifiers and text regressors Jiang et al. (2021); Desai and Durrett (2020); Glushkova et al. (2021). Other research has addressed uncertainty by prompting models to evaluate their own outputs or fine-tuning generative models to predict their own uncertainty Lin et al. (2024); Kadavath et al. (2022). However, these approaches require additional training and supervision, making them difficult to reproduce, costly to implement, and sensitive to distributional shifts.
Our work follows from a line of work inline with the concept of semantic entropy proposed in Kuhn et al. (2023); Nikitin et al. (2024); Duan et al. (2024); Lin et al. (2023). Kuhn et al. (2023) explore the entropy of the generated text by assigning semantic equivalence to the pairs of text and subsequently estimating the entropy. Similarly, Nikitin et al. (2024) and Lin et al. (2023) utilize graphical spectral analysis to enhance empirical results. However, a notable limitation in the entropy estimators proposed by Kuhn et al. (2023) and Nikitin et al. (2024) is their reliance on token likelihoods when assessing semantic equivalence, which may not always be accessible. Furthermore, Kuhn et al. (2023) acknowledge that the clustering process employed in their framework is susceptible to the order of comparisons, introducing variation into the results.
Moreover, prior work in this area has primarily focused on the empirical performance of entropy estimators. While such methods have shown promising results in practice, to the best of our knowledge, no existing studies have provided a formal theoretical analysis demonstrating that these estimators converge to the true entropy value as the sample size increases under an assumed underlying model. Investigating the theoretical properties of our proposed measure enables a deeper understanding of how factors such as the number of clusters and the sample size influence the estimator’s behavior and consistency.
Our approach aims to address these limitations by developing a robust theoretical framework for the clustering procedure, ensuring convergence properties and reducing the variability observed in prior methodologies. We introduce a theoretically analyzable metric, semantic spectral entropy, to quantify the semantic variation within a collection of texts. This uncertainty quantification measure is motivated by the observation that many generated strings, though lexically and syntactically distinct, may express equivalent semantic content. To identify such semantic equivalences, we leverage standard generative language models (LMs) as tools for semantic evaluation.
Importantly, we recognize that the language model used to assess semantic similarity is itself imperfect, introducing potential noise due to its inability to provide uniformly accurate judgments across all text pairs. To mitigate this issue, we employ spectral clustering, a well-established technique known to be statistically consistent under mild assumptions on the underlying data-generating process. This provides robustness and theoretical guarantees for our proposed measure. Specifically, we show that semantic spectral entropy is statistically consistent under a weak assumption on the language model. To the best of our knowledge, this constitutes the first semantic uncertainty measure with formally proven convergence properties.
2 Semantic spectral entropy
2.1 Semantic entropy
..We begin with a collection of $n$ textual items, denoted $\mathcal{T}=(t_{1},\ldots,t_{n})$. In contrast to the approach in Kuhn et al. (2023), we assume access only to the texts in $\mathcal{T}$, without requiring knowledge of the underlying generative process. While the methodology is designed with the goal of assessing uncertainty in outputs generated by large language models (LLMs), it is important to note that it does not necessitate $\mathcal{T}$ to be the result of a generative model. That is, the proposed method operates independently of the generation mechanism and can be applied in broader contexts. To evaluate the semantic uncertainty of the texts within a specific application domain, we introduce a theoretically grounded measure, which we term semantic spectral entropy.
While the proposed measure can be interpreted as a variance estimator for the semantics of $\mathcal{T}$, a central reason for not adopting the conventional notion of variance as a measure of uncertainty lies in the difficulty of defining a meaningful "mean" for semantic probability distributions. Unlike numerical data, semantic representations lack a natural central point around which variation can be measured. Although one might define an arbitrary reference point—such as a canonical or "standard" answer in the context of a chatbot—computing variance relative to such a reference introduces inherent bias and undermines the objectivity of the measure. Our approach, therefore, avoids reliance on predefined anchors and instead captures semantic dispersion directly, without the need for a fixed semantic center.
In contrast, entropy is a well-established measure of variation, particularly for multinomial distributions. For a distribution $\mathcal{P}(t)$ over a set of semantic clusters $\{C_{1},\cdots,C_{k}\}$, the entropy $\mathcal{E}$ is defined as:
$\mathcal{E}(t)=-\sum_{i}p(t\in C_{i})\log p(t\in C_{i}).$ (1)
This formulation captures the uncertainty or disorder associated with assigning a given text $t$ to one of the clusters. Consequently, it provides a quantitative measure of semantic uncertainty that avoids the biases introduced by arbitrary reference points.
To estimate the entropy for a given data set $t_{1},\cdots,t_{n}$, we first calculate the number of occurrences of each text $t_{i}$ in each group $C_{j}$. This is achieved by computing:
$n_{j}=\sum_{i=1}^{n}\mathbb{I}(t_{i}\in C_{j}),$
where $\mathbb{I}(t_{i}\in C_{j})$ is an indicator function that equals 1 if $t_{i}$ belongs to the cluster $C_{j}$, and 0 otherwise.
Next, the true probability $p(t\in C_{j})$ is approximated using the empirical distribution:
$\bar{p}(t\in C_{j})=\frac{n_{j}}{n},$
which represents the fraction of texts assigned to cluster $C_{j}$. Using this empirical distribution, the empirical entropy is defined as:
¯$\bar{\mathcal{E}}(\mathcal{T})=-\sum_{j}\bar{p}(t\in C_{j})\log\bar{p}(t\in C_{j}).$
This measure provides a practical estimation of semantic entropy based on observed data.
One critical step in this process is clustering the texts $t_{i}$ into disjoint groups. To do so, it is sufficient to define a relationship between $t_{i}\sim t_{j}$, such that they satisfy the properties of equivalence relation. Specifically, one needs to demonstrate
  1. 1.
    Reflexivity: For every $t_{i}$, we have $t_{i}\sim t_{i}$, meaning that any text is equivalent to itself.
  2. 2.
    Symmetry: If $t_{i}\sim t_{j}$, then $t_{j}\sim t_{i}$, meaning that equivalence is bidirectional.
  3. 3.
    Transitivity: If $t_{i}\sim t_{j}$ and $t_{j}\sim t_{k}$, then $t_{i}\sim t_{k}$, which means that equivalence is transitive.
It turns out the existence of an equivalence equation is both a necessary and sufficient condition for a definition of a breakdown of $\mathcal{T}$ into disjoint clusters Liebeck (2018). In light of this, defining $\sim$ should be based on the linguist properties of entropy measurement.
[…]
To address such challenges, Kuhn et al. (2023); Nikitin et al. (2024) propose an equivalence relation wherein $t_{i}\sim t_{j}$ if and only if $t_{i}$ is true if and only if $t_{j}$ is true. This formulation ensures that two texts, $t_{i}$ and $t_{j}$, belong to the same equivalence class if they are logically equivalent. This broader definition allows for greater flexibility and applicability in assessing semantic equivalence beyond superficial lexical similarity.Copi et al. (2016). We will present their argument as a proposition where we will put the verification in the appendix
Proposition 2.1.
The relation $t_{i}\sim t_{j}$ if "$t_{i}$ is true if and only if $t_{j}$ is true" is an equivalence relation.
In light of the fact that equivalence relations can be defined arbitrarily based on the needs of the user. We propose that the determination of equivalence relations, denoted as $\sim$, is performed through a LM that generates responses independently of the specific generation of terms $t_{1},\dots,t_{n}$. However, we do not assume that we have access to probability distribution of the tokens as proposed by Kuhn et al. (2023); Nikitin et al. (2024) which is not always available. Rather, we just require a LM which can identify this relationship. In fact, this LM can be off-the-shelf language model with a crafted prompt, something we will use in our simulation studies. The error from LM will be removed in the spectral clustering algorithm at the later stage. By leveraging this LM, we define a function $e:{\mathcal{T},\mathcal{T}}\rightarrow{0,1}$, which is formally expressed as follows:
$e(t_{i},t_{j})=\begin{cases}1&\text{if }t_{i}\sim t_{j},\\ 0&\text{otherwise.}\end{cases}$ (2)
However, since the function relies on an LM, $e(t_{i},t_{j})$ can be viewed as a Bernoulli random variable (given that we have established that there are inherent uncertainty in the LMs), whose value is dependent on the terms $t_{i}$ and $t_{j}$.Kuhn et al. (2023) did not address this issue but instead offers adopting a very powerful entailment identification model which the authors trust to identify the equivalence relation perfectly. In contrast, we suggest modeling the outputs of the LM as a random graph with an underlying distribution. In this framework, $t_{i}$ and $t_{j}$ represent nodes, while $e(t_{i},t_{j})$ are random variables that indicate the presence of an edge between the two nodes. Specifically, when $t_{i}\sim t_{j}$, the edge existence is governed by the following probability distribution:
$e(t_{i},t_{j})=\begin{cases}1&\text{with probability }p,\\ 0&\text{with probability }1-p.\end{cases}$ (3)
Conversely, when $t_{i}\not\sim t_{j}$, the edge existence follows a different probability distribution:
$e(t_{i},t_{j})=\begin{cases}1&\text{with probability }q,\\ 0&\text{with probability }1-q.\end{cases}$ (4)
To mitigate the inherent randomness introduced by the LLM, we propose leveraging spectral clustering to identify clusters of semantically similar texts.
2.2 Spectral clustering
.. To compute semantic entropy, it is crucial to identify the clusters of nodes and count the number of nodes within each cluster. Identifying these clusters in a random graph is analogous to detecting clusters in a stochastic block model Holland et al. (1983). We propose employing the spectral clustering algorithm, with the number of clusters $K$ specified in advance, as an effective approach for this task.
Spectral Clustering is a well-established algorithm for graph clustering, supported by strong theoretical foundations and efficient implementations Shi and Malik (2000); Lei and Rinaldo (2015); Su et al. (2019); Pedregosa et al. (2011). To compute semantic entropy, we aim to cluster a random graph with adjacency matrix $E$ where $E_{ij}=e(t_{i},t_{j})$, representing the pairwise similarity between text elements $t_{i}$ and $t_{j}$.
We begin by computing the Laplacian matrix $L=D-E$ where $D$ is the degree matrix. This is followed by the decomposition of the eigenvalue of $L$. Next, we construct the matrix formed by the first $K$ eigenvectors of $L$ denoted $\hat{U}\in\mathbb{R}^{n\times K}$. This matrix serves as input to an appropriate $(1+\epsilon)-$ k-means clustering algorithm Kumar et al. (2004); Choo et al. (2020).
The output of this procedure is $K$ distinct clusters $C_{1},\cdots C_{K}$. For each text element $t_{i}$, we assign a corresponding vector $g_{i}$ where
$g_{ij}=\begin{cases}1\text{ if }t_{i}\in C_{j}\\ 0\text{ otherwise }\end{cases}$
This binary indicator vector $g_{i}$ encodes the cluster membership for each text element $t_{i}$
Finally, we compute the estimated entropy based on the number of texts within each cluster. The entropy $\hat{\mathcal{E}}$ can be approximated using the following formula:
^$\hat{\mathcal{E}}(\mathcal{T})=-\sum_{j=1}^{k}\hat{p}(C_{j})\log(\hat{p}(C_{j})),$ (5)
where $\hat{p}(C_{j})=\frac{1}{n}\sum_{i=1}^{n}g_{ij}$.This expression represents the empirical entropy based on the distribution of texts among the $K$ clusters, providing a measure of the uncertainty or diversity within the semantic structure of the data.
2.3 Full algorithm and implementation
.. We merge the process of finding sermantic entropy with spectral clustering to present the full algorithm as Algorithm 1: Sermantic Spectral Entropy.
  Begin with $\mathcal{T}=\{t_{1},\cdots t_{n}\}$
  for $i,j\in\{1,\cdots n\}\times\{1,\cdots n\},i\neq j$ do
     Use LLM to compute $E_{i,j}=e(t_{i},t_{j})$.
  end for
  Find the Laplacian of $E$, $L=D-E$
  Compute the first $K$ eigenvectors $u_{1},\dots,u_{k}$ of $L$ and the top $K$ eigenvalues $\lambda_{1},\cdots\lambda_{k}$.
  Let $\hat{U}\in\mathbb{R}^{n\times k}$ be the matrix containing the vectors $u_{1},\dots,u_{k}$ as columns.
  Use $(1+\epsilon)$ K-means clustering algorithm to cluster the rows of $U$
  Let $g_{ij}$ be an $(1+\epsilon)-$approximate solution to a $K-$means clustering algorithm
  Compute ^$\hat{\mathcal{E}}(\mathcal{T})$ using $g_{ij}$
Algorithm 1 Sermantic Spectral Entropy
This polynomial-time algorithm is characterized by the largest computational cost associated with the determination of $E_{ij}$. However, computing $E_{ij}$ is embarrassingly parallel, meaning that it can be efficiently distributed across multiple processing units. Furthermore, there are well-established implementation, such as Microsoft Azure’s Prompt-Flow Esposito (2024) and LangChain Mavroudis (2024) that facilitate the implementation of parallel workflows, making it feasible to deploy such parallelized tasks with relative ease.
2.4 Finding K
.. A notable limitation of this analysis is the unavailability of $K$ in the direct computation of semantic spectral entropy. However, the determination of $K$ for stochastic block model has been well studied Lei (2016); Wang and Bickel (2017); Chen and Lei (2018). We will describe the cross-validation approach Chen and Lei (2018) in detail. The principle behind cross-validation involves predicting the probabilities associated with inter-group connections ($p$) and intra-group connections ($q$). If the estimated value of $K$ is too small, it fails to accurately recover the true underlying probabilities; conversely, if $K$ is too large, it leads to overfitting to noisy data. This approach has the potential to recover the true cluster size under relatively mild conditions.
3 Theoretical Results
.. Our theoretical analysis involves a proof that the estimator is strongly consistent, i.e. the estimator converges to true value almost surely, and an analysis of its rate with respect to the number of cluster $K$.
We divide our analysis into two subsections. The first subsection examines a fixed set of $\mathcal{T}={t_{1},\dots,t_{n}}$, which is assumed to exhibit some inherent clusters $C_{1},\dots,C_{K}$. Under the assumption of perfect knowledge of these clusters, the empirical entropy $\bar{\mathcal{E}}$ can be determined. The primary focus in this subsection is on the performance of spectral clustering algorithms. The second subsection explores a scenario in which there exists an underlying generative mechanism that allows for the infinite generation of $t_{i}$. In this case, we permit $K$ to increase with $n$, though at a significantly slower rate. This scenario is particularly relevant for evaluating the performance of RAG in the context of continuous generation of results in response to a given query.
3.1 Performance of spectral clustering algorithms
[…]
Lemma 3.1.
Suppose that there exists $10<c_{2}<$ 1 such that $2Kn_{\min}/n\geq c_{2}$,
|^$|\hat{\mathcal{E}}(\mathcal{T})-\bar{\mathcal{E}}(\mathcal{T})|\leq h\left(\frac{2K}{c_{2}}\right)\left|\frac{1}{n}(M_{\text{error}})\right|$ (6)
where $h(x)=\left(x+\log\left(x\right)\right)$.
The proof is presented in the Appendix section B.2.1. We begin by presenting the result of strong consistency for the spectral clustering algorithm.
Theorem 3.2.
Under regularity conditions, the estimated entropy empirical entropy ^$\hat{\mathcal{E}}(\mathcal{T})$ is strongly consistent with the empirical entropy, i.e.
[…]
The proof is provided in the Appendix section B.2. This establishes strong consistency result that we aim to present. At the same time, we also want to show the finite sample properties of the estimator ^$\hat{\mathcal{E}}(\mathcal{T})$.
Theorem 3.3.
If there exists $0<c_{2}\leq 1$ and $\lambda>0$ such that $2Kn_{\min}/n\geq c_{2}$, and $p=\alpha_{n}=\alpha_{n}(q+\lambda)$, where $\alpha_{n}\geq\log(n)$ then with probability at least $1-\frac{1}{n}$
|¯$|\bar{\mathcal{E}}(\mathcal{T})-\hat{\mathcal{E}}(\mathcal{T})|\leq h\left(\frac{2K}{c_{2}}\right)\frac{n_{\max}}{4c_{2}^{2}n_{\min}^{2}\alpha_{n}K^{2}}$ (8)
where $h(x)=\left(x+\log\left(x\right)\right)$, $n_{\max}=\max_{j}\{n_{j}:j=1,\dots K\}$, and $n_{\min}=\min_{j}\{n_{j}:j=1,\dots K\}$.
The full proof is provided in the appendix section B.3. A brief outline of the proof is as follows: we begin by using the results from Lei and Rinaldo (2015), which establish the rate of convergence for the stochastic block model. Next, we relate the errors of the spectral clustering algorithm to the errors in the empirical entropy, using the lemma 3.1 to establish this connection.
Remark.
This result is particularly relevant for computing semantic entropy, as the output generated by LMs is produced with a probability that is independent of $n$. As a result, we have $\alpha_{n}=O(1)$. Assuming balanced community sizes, the convergence rate is therefore $O(\frac{1}{n})$. This is formally stated in the following corollary:
Corollary 3.3.1.
If there exists a constant $0<c_{2}\leq 1$ such that $2Kn_{\min}/n\geq c_{2}$ and $\alpha_{n}=alpha>0$, then there exists a constant $\alpha$ such that with probability at least $1-\frac{1}{n}$,
|¯$|\bar{\mathcal{E}}(\mathcal{T})-\hat{\mathcal{E}}(\mathcal{T})|\leq h\left(\frac{2K}{c_{2}}\right)\frac{1}{c_{2}^{4}\alpha n}.$ (9)
The proof of this result is provided in the Appendix section B.3.1.
Remark.
In particular, we observe that the convergence rate is $O\left(\frac{1}{n}\right)$. This means that the error associated with spectral clustering is small, and our estimated entropy converges to the empirically entropy quickly.
3.2 Performance under a generative model
.. In practical terms, we assume the presence of a generator, specifically an RAG, that produces identically distributed independent random variables $t_{i}$’ that collectively form semantic clusters $C_{1}\dots C_{K}$. In essence, we have $t_{i}\sim G$ such that $t_{i}\in C_{j}$ with probability $p(C_{j})$. In this model, there is a true value of entropy $\mathcal{E}(\mathcal{T})$ given in Equation 1, and we want to find the convergence rate of our method.
Theorem 3.4.
If there exists a constant $\alpha$ such that $p=\alpha=\alpha(q+\lambda)$, then with probability at least $1-\frac{3}{n}$,
|^$\begin{array}[]{cc}|\mathcal{E}-\hat{\mathcal{E}}|&\leq h\left(\frac{1}{p_{\min}}\right)K\sqrt{\frac{1}{2n}\log\left(2Kn\right)}\\ &+h\left(\frac{1}{m(n)p_{\min}}\right)\frac{1}{16K^{4}m(n)^{4}p_{\min}^{4}n}\end{array}$ (10)
where $m(n)=\left(1-\sqrt{2\log(nK)/np_{\min}}\right)$ and $p_{\min}=\min\{p(C_{1})\dots p(C_{K})\}$.
Most of the material used for this proof is presented in Corollary 3.3.1.
Proof.
Consider the following equality
$|\mathcal{E}-\hat{\mathcal{E}}|\leq|\mathcal{E}-\bar{\mathcal{E}}+\bar{\mathcal{E}}-\hat{\mathcal{E}}|\leq|\mathcal{E}-\bar{\mathcal{E}}|+|\bar{\mathcal{E}}-\hat{\mathcal{E}}|,$
We know that there are three sufficient conditions for Equation 10. These are
  1. C1:
    |¯$|\mathcal{E}-\bar{\mathcal{E}}|\leq h\left(\frac{1}{p_{\min}}\right)K\sqrt{\frac{1}{2n}\log\left(2Kn\right)},$
  2. C2:
    $\exists c_{2}\text{ such that }0<c_{2}\leq 1$ and $2Kn_{\min}/n\geq c_{2},$
  3. C3:
    |¯$|\bar{\mathcal{E}}(\mathcal{T})-\hat{\mathcal{E}}(\mathcal{T})|\leq h\left(\frac{2K}{c_{2}}\right)\frac{1}{c_{2}^{4}n}.$
Then, using union bound
[…]
In Lemma B.6 and B.7 of the appendix, we show that |¯$|\mathcal{E}-\bar{\mathcal{E}}|\geq h\left(\frac{1}{p_{\min}}\right)K\sqrt{\frac{1}{2n}\log\left(2Kn\right)}$ with probability at most $\frac{1}{n}$.
In Lemma B.8 of the Appendix, we show that setting $c_{2}=2K\left(1-\sqrt{\frac{2\log(nK)}{np_{\min}}}\right)p_{\min}$, we have $2Kn_{\min}/n<c_{2}$ with probability at most $\frac{1}{n}$.
Finally, the corollary 3.3.1 tells us that |¯$|\bar{\mathcal{E}}(\mathcal{T})-\hat{\mathcal{E}}(\mathcal{T})|>h\left(\frac{2K}{c_{2}}\right)\frac{1}{c_{2}^{4}n}$ occurs with probability at most $\frac{1}{n}$. ∎
Remark.
One observation is that the empirical entropy converges to true entropy at a rate slower than that of estimated entropy to the empirical entropy. This is natural since each $t_{i}$ has the opportunity to make a $n-1$ connection with other $t_{j}$s, resulting in $n(n-1)/2$ independent observations, whereas each generator generates only $n$ independent observations.
3.3 Discussion on $K$
.. An intriguing question to consider is the rate at which $K$, the number of clusters, can grow with $n$, the number of texts, as it is natural to expect $K$ to increase with $n$. Focusing solely on the spectral clustering algorithm, the error is characterized as $O((K+\log(K))/n)$. Thus, under the condition $K=o(n^{1-\delta})$ for some $\delta>0$, we have |¯$|\bar{\mathcal{E}}(\mathcal{T})-\hat{\mathcal{E}}(\mathcal{T})|\to 0$ in probability. In contrast, when considering a scenario involving a generative model, a stricter condition is required. Specifically, $K$ must satisfy $K=o(n^{1/2-\delta})$, with $\delta>0$, to ensure $|\mathcal{E}(\mathcal{T})-\hat{\mathcal{E}}(\mathcal{T})|\to 0$ in probability.
4 Simulation and data studies
.. As this paper focuses more on the theoretical analysis of semantic spectral entropy with respect to variable $n$ and $K$, we decide against using the evaluation method proposed in Kuhn et al. (2023); Duan et al. (2024); Lin et al. (2023) in favor of constructing a simulation where we know the true entropy $\bar{\mathcal{E}}$. This allows us to better analyze how $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ changes with choice of generator $e$, $K$ and $n_{\min}$.
To construct a non-trivial simulation for this use case, we evaluate the performance of our algorithms within the context of an unordered set of elementary proposition statements that has no logical interconnections. This approach draws upon the philosophical framework defined by (Wittgenstein, 2023) in Tractatus Logico-Philosophicus, where each elementary proposition represents a singular atomic fact. Within this framework, texts containing an identical set of elementary propositions are deemed semantically equivalent. The primary advantage of this experimental design lies in its efficiency, as it facilitates the generation of thousands of samples with minimal generator propositions, all while maintaining knowledge of the ground truth.
For example, we can consider a list of things that a hypothetical individual "John" likes to do in his free time:
  • Running/Jogging
To generate a cluster of text from this set of hobbies, we begin by randomly selecting $M$ items from a total of $N$ items in the list to formulate the compound proportion. This selection process yields $\binom{N}{M}$ potential subset of hobbies and we know that two subsets of hobbies are the same as long as their elements are the same. Next, to create individual text samples $t_{i}$ within the group, we randomly permute the order of the $M$ selected elements in the subset. This permutation process generates $M!$ unique samples for each combination of hobbies. Finally, the hobbies are placed in their permuted order in a sentence like that in the following.
"In his free time, John likes hobby $1$, hobby $2$, hobby $3$, …, and hobby $M$ as his hobbies."
In order to prevent models from relying on sentence structure, some of these sentences are being designed.
We replicate this simulation setup in two different settings. In the first setting (Experiment 1), we consider ten common hobbies that a hypothetical individual engages in during their spare time. In the second setting (Experiment 2), we examine ten historical events that occurred on December 3, which we collect from Wikipedia Wikipedia contributors (2024). These two settings allow us to analyze the impact of text length by comparing long and short text scenarios.
We utilize Microsoft Phi-3.5 Abdin et al. (2024), OpenAI GPT3.5-turbo (denoted as GPT) Hurst et al. (2024), A21-Jamba 1.5 Mini (denoted as A21) Lieber et al. (2021), Cohere-command-r-08-2024 (denoted as Cohere) Ustun et al. (2024), Ministral-3B (denoted as Minstral) Jiang et al. (2023) and the Llama 3.2 70B model (denoted as Llama) Dubey et al. (2024) as $e$. These models are lightweight, off-the-shelf language models that are cost-effective to deploy and exhibit efficiency in generating outputs, thereby offsetting the computational cost of determining sermantic relationships. The exact prompt used to generate the verdict is specified in Appendix D.  [^1] Code are in https://github.com/yiliu9090/sermantic-spectral-entropy
Experiment 1 (Hobbies)
ratio 0.2,0.3,0.5 0.3,0.3,0.4 0.5,0.5 $p-q$ $p$ q
datasize 30 50 70 30 50 70 30 50 70
LLAMA 0.36 0.49 0.44 0.34 0.43 0.46 0.30 0.27 0.26 0.17 0.17 0.00
MINISTRAL 0.22 0.27 0.13 0.25 0.23 0.21 0.14 0.22 0.21 0.22 0.99 0.77
COHERE 0.04 0.02 0.06 0.02 0.03 0.00 0.00 0.00 0.00 0.55 0.61 0.05
A21 0.05 0.00 0.00 0.00 0.01 0.00 0.00 0.00 0.00 0.81 0.96 0.15
PHI 0.08 0.07 0.07 0.03 0.03 0.00 0.00 0.00 0.00 0.67 0.67 0.01
GPT 0.06 0.02 0.00 0.01 0.00 0.00 0.00 0.00 0.00 0.80 0.87 0.07
Experiment 2 (Historical Events)
LLAMA 0.04 0.02 0.05 0.03 0.03 0.07 0.04 0.00 0.00 0.75 0.75 0.00
MINISTRAL 0.20 0.20 0.19 0.17 0.24 0.19 0.04 0.07 0.19 0.25 1.00 0.75
COHERE 0.08 0.04 0.09 0.08 0.09 0.04 0.00 0.01 0.04 0.52 0.98 0.46
A21 0.09 0.14 0.12 0.08 0.00 0.01 0.00 0.00 0.00 0.49 0.50 0.01
PHI 0.10 0.04 0.19 0.16 0.03 0.06 0.07 0.06 0.00 0.28 0.28 0.00
GPT 0.02 0.01 0.06 0.02 0.04 0.00 0.00 0.00 0.00 0.76 0.96 0.20
Table 1: Average $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ over simulation 10 iterations. We have three different ratio value run over three different data sizes. For $e$, we use Microsoft Phi-3.5 Abdin et al. (2024), OpenAI GPT3.5-turbo Hurst et al. (2024), A21-Jamba 1.5 Mini Lieber et al. (2021), Cohere-command-r-08-2024 Ustun et al. (2024), Ministral-3B Jiang et al. (2023) and the Llama 3.2 70B model Dubey et al. (2024). We also present the average $p$, $q$ and $p-q$ observed in the experiments.
Figure 1: A scatter plot of $p-q$ against $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ for outcome in experiment 1. The different colors represents different language models used as $e$: A21 in blue, Phi in Orange, GPT in Green, Cohere in Red, Llama is Purple and Ministral in Brown. We notice that there is clear phrase change point where for $p-q<0.4$, we have that $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ is very high most of the time, for $p-q>0.4$, $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ is small with occasional jumps that the theory predicts.
We conduct simulation studies for ratio configurations of (0.2, 0.3, 0.5), (0.3, 0.3, 0.4), and (0.5, 0.5), with sample sizes of 30, 50, and 70. The average deviation, $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$, over 10 iterations using different models as $e$ is recorded in Table 1.
In Experiment 1, the algorithm performs strongly when using Cohere, A21, Phi, and GPT, whereas its performance is weaker with Minstral and Llama. In Experiment 2, Llama, Cohere, A21, and GPT demonstrate strong performance, while Minstral and Phi exhibit weaker results. These findings suggest that Llama performs more effectively with longer text, which we primarily attribute to differences in $p-q$. Specifically, in Experiment 1, $p-q$ is small for Llama and Minstral but large for Cohere, A21, Phi, and GPT. In Experiment 2, $p-q$ is small for Phi and Minstral but large for Llama, A21, Cohere, and GPT.
A comparative analysis of Llama and Minstral in Experiment 1, as well as Phi and Minstral in Experiment 2, indicates that neither $p$ nor $q$ alone has a strong effect on $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$. Instead, when plotting $p-q$ against $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ in Figure 1, we observe a phase change at $p-q=0.4$. Specifically, for $p-q<0.4$, $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ is relatively high, whereas for $p-q>0.4$, $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ tends to be lower. Notably, the experimental setup—defined by the number of data points and cluster ratio—has a considerably weaker effect on $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$.
5 Discussion
.. Many natural language processing tasks exhibit a fundamental invariance: sequences of distinct tokens can convey identical meanings. This paper introduces a theoretically grounded metric for quantifying semantic uncertainty, referred to as semantic spectral clustering. This approach reframes the challenge of measuring semantic uncertainty as a prompt-engineering problem, which can be applied to any large language model (LLM), as demonstrated through our simulation analysis. In addition, unsupervised uncertainty can offer a solution to the issue identified in prior research, where supervised uncertainty measures face challenges in handling distributional shifts. While we define two texts as having equivalent meaning if and only if they mutually imply one another, alternative definitions may be appropriate for specific use cases. For example, legal documents could be clustered based on the adoption of similar legal strategies, with documents grouped together if they demonstrate comparable approaches. In such scenarios, the entropy of the legal documents could also be computed to quantify their informational diversity. We have demonstrated that, provided there exists a function $e$ capable of performing the evaluation with weak accuracy, this estimator remains consistent. Given the reasoning capabilities of large language models (LLMs), we foresee numerous possibilities for extending this method to a wide range of applications.
In addition to the methodology presented, we present a theoretical analysis of the proposed algorithms by proving a theorem concerning the contraction rates of the entropy estimator and its strong consistency. Although the algorithm utilizes generative models, which are typically treated as black-boxes, we simplify the analysis by considering the outputs of these models as random variables. We demonstrate that only a few conditions on the generative are sufficient for our spectral clustering algorithm to achieve strong consistency. Our approach allows for many statistical methodologies to be applied in conjunctions with generative models to analyze text at a level previously not achievable by humans.
Appendix A Limitation
.. We acknowledge that, while this research offers a theoretically consistent measurement of variation, it does not account for situations where two pieces of text may partially agree. For example, two texts may contain points of agreement and points of disagreement. This is particularly common when different authors cite similar sources, but reach contradictory conclusions. Computing semantic similarity in this case is difficult.
Appendix B Theoretical Result
B.1 Proof of proposition 2.1
Proof.
To prove that the relation $t_{i}\sim t_{j}$ if $t_{i}$ is true if and only if $t_{j}$ is true is an equivalence relation, we need to meet 3 key criteria, namely symmetry, reflexivity, and Transitivity.
First, symmetry $t_{i}\sim t_{j}$ implies that $t_{j}$ is true $\Leftrightarrow$ $t_{j}$ is true, but this also means $t_{j}$ is true $\Leftrightarrow$ $t_{i}$ is true. Then we have $t_{j}\sim t_{i}$.
Second, reflexivity, $t_{i}\sim t_{j}$ implies $t_{j}$ is true $\Leftrightarrow$ $t_{j}$ is true. But this means that $t_{j}$ is true $\Leftrightarrow$ $t_{i}$ is true. Then we have $t_{i}\sim t_{j}$.
Third, transitivity, If $t_{i}\sim t_{j}$ and $t_{j}\sim t_{k}$, Then if $t_{i}$ is true $\Rightarrow$ $t_{j}$ is true $\Rightarrow$ $t_{k}$ is true, which means $t_{i}$ is true $\Rightarrow$ $t_{k}$ is true. On the other hand, using the same argument, $t_{k}$ is true $\Rightarrow$ $t_{j}$ is true $\Rightarrow$ $t_{i}$ is true. This means the $t_{k}$ is true $\Rightarrow$ $t_{i}$ is true. Therefore $t_{i}\sim t_{k}$.
The three points is sufficient to demonstrate that $\sim$ is a equivalence relation. ∎
B.2 Proof of Theorem 3.2
To prove Theorem 3.2, we adopt notations from Su et al. (2019). Consider the adjacency matrix $E$ which is determined by a Language model.
Let $d_{i}=\sum_{j=1}^{n}E_{ij}$ denote the degree of node $i$, $D=\text{diag}(d_{1},\cdots,d_{n})$, and $L=D^{-1/2}ED^{-1/2}$ be the graph Laplacian. We also define $n_{k}$ be the number of text in each cluster. We denote a block probability matrix $B=B_{k_{1}k_{2}}$ where $k_{1},k_{2}\in\{1,\cdots K\}$ be the clusters index. i.e.
$B_{k_{1}k_{2}}=\begin{cases}p\quad\text{if $k_{1}=k_{2}$}\\ 1-q\quad\text{otherwise.}\end{cases}$
Let $\mathbb{E}(E)=P$ i.e. the probability of edge between $i$ and $j$ is given by $P_{ij}=B_{k_{1}k_{2}}$ if text $i$ is in $C_{k_{1}}$ and $j$ is in $C_{k_{2}}$. Denote $Z=\{Z_{ik}\}$ be a $n\times K$ binary matrix providing the cluster membership of text $t$, i.e., $Z_{ik}=1$ if text $i$ is in $C_{k}$ and $Z_{ik}=0$ otherwise. The population version of the Laplacian is given by $\mathcal{L}=\mathcal{D}^{-1/2}P\mathcal{D}^{-1/2}$ where $\mathcal{D}=\text{diag}(d_{1}\cdots d_{n})$ where $d_{i}=\sum_{j=1}^{n}P_{ij}=p+(n-1)(q)$.
Let $\pi_{kn}=n_{k}/n,W_{k}=\sum_{l=1}^{K}B_{kl}\pi_{ln}$, $\mathcal{D}_{B}=\text{diag}(W_{1},\cdots W_{K})$, and $B_{0}=\mathcal{D}_{B}^{-1/2}B\mathcal{D}_{B}^{-1/2}$
Assumption B.1 (Assumption 1 in Su et al. (2019)).
$P$ is rank $k$ and spectral decomposition $\Pi_{n}^{1/2}P\Pi_{n}^{1/2}$ is $S_{n}\Omega_{n}S_{n}^{T}$ in which $S_{n}$ is a $K\times K$ matrix such that $S_{n}^{T}S_{n}=I_{K\times K}$ and $\Omega_{n}=\text{diag}(\omega_{1}\cdots\omega_{K_{n}})$ such that $|\omega_{1}|\geq|\omega_{2}|\geq\cdots\geq|\omega_{K_{n}}|$
Assumption B.1 implies that the spectral decomposition
$\mathcal{L}=U_{n}\Sigma_{n}U_{n}^{T}=U_{1n}\Sigma_{1n}U_{1n}^{T}$
where $\Sigma_{n}=\operatorname{diag}\left(\sigma_{1n},\ldots,\sigma_{Kn},0,\ldots,0\right)$ is a $n\times n$ matrix that contains the eigenvalues of $\mathcal{L}$ such that $\left|\sigma_{1n}\right|\geq\left|\sigma_{2n}\right|\geq\cdots\geq\left|\sigma_{Kn}\right|>0,\Sigma_{1n}=\operatorname{diag}\left(\sigma_{1n},\ldots,\sigma_{Kn}\right)$, the columns of $U_{n}$ contain the eigenvectors of $\mathcal{L}$ associated with the eigenvalues in $\Sigma_{n},U_{n}=\left(U_{1n},U_{2n}\right)$, and $U_{n}^{T}U_{n}=I_{n}$ Su et al. (2019).
Assumption B.2 (Assumption 2 in Su et al. (2019)).
There exists constant $C_{1}>0$ and $c_{2}>0$ such that
$C_{1}\geq\lim\sup_{n}\sup_{k}n_{k}K/n\geq\lim\inf_{n}\inf_{k}n_{k}K/n\geq c_{2}$
Assumption B.3 (Assumption 3 in Su et al. (2019)).
Let $\mu_{n}=\min_{i}d_{i}$ and $\rho_{n}=\max(\sup_{k_{1}k_{2}}[B_{0}]_{k_{1}k_{2}},1)$. Then $n$ sufficiently large,
$\frac{K\rho_{n}\log^{1/2}(n)}{\mu_{n}^{1/2}\sigma_{Kn}^{2}}\left(1+\rho_{n}+\left(\frac{1}{K}+\frac{\log(5)}{\log(n)}\right)^{1/2}\rho_{n}^{1/2}\right)\leq 10^{-8}C_{1}^{-1}c_{2}^{1/2}.$
Let
$\hat{O}_{n}=\bar{U}\bar{V}^{T}$
where $\bar{U}\bar{\Sigma}\bar{V}^{T}$ is the singular value decomposition of $\hat{U}_{1n}^{T}U_{1n}$. we also denote $\hat{u}_{1i}^{T}$ and $u_{1i}^{T}$ as the $i$-th rows of $\hat{U}_{1n}$ and $U_{1n}$, respectively.
Now we present the notation of the K-means algorithm. With a little abuse of notation, let $\hat{\beta}_{\text{in }}\in\mathbb{R}^{K}$ be a generic estimator of $\beta_{g_{i}^{0}n}\in\mathbb{R}^{K}$ for $i=1,\ldots,n$. To recover the community membership structure (i.e., to estimate $g_{i}^{0}$ ), it is natural to apply the K-means clustering algorithm to $\left\{\widehat{\beta}_{\text{in }}\right\}$. Specifically, let $\mathcal{A}=\left\{\alpha_{1},\ldots,\alpha_{K}\right\}$ be a set of $K$ arbitrary $K\times 1$ vectors: $\alpha_{1},\ldots,\alpha_{K}$. Define
$\widehat{Q}_{n}(\mathcal{A})=\frac{1}{n}\sum_{i=1}^{n}\min_{1\leq l\leq K}\left\|\hat{\beta}_{in}-\alpha_{l}\right\|^{2}$
and $\widehat{\mathcal{A}}_{n}=\left\{\widehat{\alpha}_{1},\ldots,\widehat{\alpha}_{K}\right\}$, where $\widehat{\mathcal{A}}_{n}=\arg\min_{\mathcal{A}}\widehat{Q}_{n}(\mathcal{A})$. Then we compute the estimated cluster identity as
$\hat{g}_{i}=\underset{1\leq l\leq K}{\arg\min}\left\|\hat{\beta}_{\text{in }}-\widehat{\alpha}_{l}\right\|,$
where if there are multiple $l$ ’s that achieve the minimum, $\hat{g}_{i}$ takes value of the smallest one. We then state the key assumption that relates to K-means clustering algorithm.
Assumption B.4 (Assumption 7 in Su et al. (2019)).
Suppose for $n$ sufficiently large,
$15C^{*}\frac{K\rho_{n}\log^{1/2}(n)}{\mu_{n}^{1/2}\sigma_{Kn}^{2}}\left(1+\rho_{n}+\left(\frac{1}{K}+\frac{\log(5)}{\log(n)}\right)^{1/2}\rho_{n}^{1/2}\right)\leq c_{2}C_{1}^{-1/2}\sqrt{2}$
Where $C^{*}=3528C_{1}c_{2}^{-1/2}$
Theorem B.1.
(Collorary 2.2) Corollary 2.2. Suppose that Assumptions B.1, B.2, B.3, and B.4 hold and the $K$-means algorithm is applied to $\hat{\beta}_{in}=(n/K)^{1/2}\hat{u}_{1i}$ and $\beta_{g_{i}^{0}n}=(n/K)^{1/2}\hat{O}_{n}u_{1i}$ Then,
[…]
We now have define the error of mis-classification.
Definition B.1.
Denote $M_{\text{error}}=\sum_{j}\sum_{i}\mathbb{I}(g_{ij}\neq g^{\text{True}}_{ij})$ as the mis-classification error.
Lemma B.2.
[…]
Proof.
Notice $\mathbb{I}(g_{ij}\neq g^{\text{True}}_{ij})$ can only takes up value $1$ or 00. Therefore $\sum_{j}\sum_{i}\mathbb{I}(g_{ij}\neq g^{\text{True}}_{ij})\neq 0\Leftrightarrow\exists i,j\text{ s.t }\mathbb{I}(g_{ij}\neq g^{\text{True}}_{ij})\neq 0\Leftrightarrow\sup_{i,j}\mathbb{I}(g_{ij}\neq g^{\text{True}}_{ij})\neq 0$
[…]
Here we use the classical notation i.o. as happens infinitely often. ∎
Lemma B.3.
$\sum_{j}\left|\sum_{i=1}^{n}g_{ij}-n_{j}\right|\leq M_{\text{error}}$
Proof.
$\sum_{j}\left|\sum_{i=1}^{n}g_{ij}-n_{j}\right|$ $=\sum_{j}\Biggl{|}\sum_{i}\mathbb{I}(g_{ij}=1,g^{\text{True}}_{ij}=0)+\mathbb{I}(g_{ij}=1,g^{\text{True}}_{ij}=1)+\mathbb{I}(g_{ij}=0,g^{\text{True}}_{ij}=1)$
$-\mathbb{I}(g_{ij}=0,g^{\text{True}}_{ij}=1)-n_{j}\Biggr{|}$
$=\sum_{j}\Biggl{|}\sum_{i}\mathbb{I}(g_{ij}=1,g^{\text{True}}_{ij}=0)-\mathbb{I}(g_{ij}=0,g^{\text{True}}_{ij}=1)$
$+\sum_{i}\mathbb{I}(g_{ij}=0,g^{\text{True}}_{ij}=1)+\mathbb{I}(g_{ij}=0,g^{\text{True}}_{ij}=1)-n_{j}\Biggr{|}$
$=\sum_{j}\Biggl{|}\sum_{i}\mathbb{I}(g_{ij}=1,g^{\text{True}}_{ij}=0)-\mathbb{I}(g_{ij}=0,g^{\text{True}}_{ij}=1)+n_{j}-n_{j}\Biggr{|}$
$=\sum_{j}\Biggl{|}\sum_{i}\mathbb{I}(g_{ij}=1,g^{\text{True}}_{ij}=0)-\mathbb{I}(g_{ij}=0,g^{\text{True}}_{ij}=1)\Biggr{|}$
$\leq\sum_{j}\sum_{i}\mathbb{I}(g_{ij}=1,g^{\text{True}}_{ij}=0)+\mathbb{I}(g_{ij}=0,g^{\text{True}}_{ij}=1)$
$=\sum_{j}\sum_{i}\mathbb{I}(g_{ij}\neq g^{\text{True}}_{ij})$
$=M_{\text{error}}$
B.2.1 Proof of lemma 3.1
Now we prove lemma 3.1.
Proof.
Recall that
  • $\hat{p}(C_{j})=\frac{1}{n}\sum_{i=1}^{n}g_{ij}$ and ^$\hat{\mathcal{E}}(\mathcal{T})=-\sum_{j=1}^{K}\hat{p}(C_{j})\log(\hat{p}(C_{j}))$
  • $\bar{p}(C_{j})=\frac{n_{j}}{n}$ and ¯$\bar{\mathcal{E}}(\mathcal{T})=-\sum_{j=1}^{K}\bar{p}(C_{j})\log(\bar{p}(C_{j}))$
[…]
$=\left|\sum_{j=1}^{K}\hat{p}(C_{j})\log(\hat{p}(C_{j}))-\hat{p}(C_{j})\log(\bar{p}(C_{j}))+\hat{p}(C_{j})\log(\bar{p}(C_{j}))-\bar{p}(C_{j})\log(\bar{p}(C_{j}))\right|$
$=\left|\sum_{j=1}^{K}\hat{p}(C_{j})\log\left(\frac{\hat{p}(C_{j})}{\bar{p}(C_{j})}\right)-\left(\hat{p}(C_{j})-\bar{p}(C_{j})\right)\log(\bar{p}(C_{j}))\right|$
$=\left|\sum_{j=1}^{K}\hat{p}(C_{j})\log\left(\frac{\hat{p}(C_{j})}{p(C_{j})}\right)-\left(\hat{p}(C_{j})-\bar{p}(C_{j})\right)\log(\bar{p}(C_{j}))\right|$
$\leq\left|\sum_{j=1}^{K}\hat{p}(C_{j})\log\left(\frac{\hat{p}(C_{j})}{\bar{p}(C_{j})}\right)\right|+\left|\sum_{j=1}^{K}\left(\hat{p}(C_{j})-\bar{p}(C_{j})\right)\log(\bar{p}(C_{j}))\right|$
$\leq\left|\sum_{j=1}^{K}\left(\frac{\hat{p}(C_{j})-\bar{p}(C_{j})}{\bar{p}(C_{j})}\right)\right|+\left|\sum_{j=1}^{K}\left(\hat{p}(C_{j})-\bar{p}(C_{j})\right)\log(\bar{p}(C_{j}))\right|$
$=\left|\sum_{j=1}^{K}\left(\frac{\frac{1}{n}\sum_{i=1}^{n}g_{ij}-\bar{p}(C_{j})}{\bar{p}(C_{j})}\right)\right|+\left|\sum_{j=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}g_{ij}-\bar{p}(C_{j})\right)\log(\bar{p}(C_{j}))\right|$
$=\left|\sum_{j=1}^{K}\left(\frac{\frac{1}{n}\left(\sum_{i=1}^{n}g_{ij}-n_{j}\right)}{\bar{p}(C_{j})}\right)\right|+\left|\sum_{j=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}g_{ij}-\bar{p}(C_{j})\right)\log(\bar{p}(C_{j}))\right|$
$\leq\sum_{j=1}^{K}\left|\frac{\frac{1}{n}\left(\sum_{i=1}^{n}g_{ij}-n_{j}\right)}{\bar{p}(C_{j})}\right|+\sum_{j=1}^{K}\left|\frac{1}{n}\sum_{i=1}^{n}(g_{ij}-n_{j})\right|\left|\log(\bar{p}(C_{j}))\right|$
$\leq\left|\frac{\frac{2K}{n}(M_{\text{error}})}{c_{2}}\right|+\log\left(\frac{2K}{c_{2}}\right)\left|\frac{1}{n}(M_{\text{error}})\right|$
$=h\left(\frac{2K}{c_{2}}\right)\left|\frac{1}{n}(M_{\text{error}})\right|$
where $h(x)=\left(x+\log\left(x\right)\right)$. ∎
We prove Theorem 3.2. To do so, we first restate Theorem 3.2 with all the conditions required to get to the outcome.
Theorem B.4 (Theorem 3.2 with all conditions stated).
Assume that Assumptions B.1, B.2, B.3, and B.4 hold and the $K$-means algorithm is applied to $\hat{\beta}_{in}=(n/K)^{1/2}\hat{u}_{1i}$ and $\beta_{g_{i}^{0}n}=(n/K)^{1/2}\hat{O}_{n}u_{1i}$ Then
[…]
Proof.
Using Theorem B.1, we know that under Assumptions B.1, B.2, B.3, and B.4, we have that
[…]
Using Lemma B.2, we know that
[…]
Using results from Lemma 3.1, we know that $M_{\text{error}}\rightarrow 0\quad a.s.\Rightarrow\hat{\mathcal{E}}(\mathcal{T})\rightarrow\bar{\mathcal{E}}(\mathcal{T})\quad a.s.$. ∎
Now we try to prove Theorem 3.3. To do so, we state corollary 3.2 in Lei and Rinaldo (2015).
B.3 Proof of Theorem 3.3
Theorem B.5 (Corollary 3.2 in Lei and Rinaldo (2015)).
Let $E$ be an adjacency matrix from the $\operatorname{SBM}(Z,B)$, where $B=\alpha_{n}B_{0}$ for some $\alpha_{n}\geq\log n/n$ and with $B_{0}$ having minimum absolute eigenvalue $\geq\lambda>0$ and $\max_{k\ell}B_{0}(k,\ell)=1$. Let $g_{ij}$ be the output of spectral clustering using $(1+\varepsilon)$-approximate $k$-means. Then there exists an absolute constant $c$ such that if
$(2+\varepsilon)\frac{Kn}{n_{\min}^{2}\lambda^{2}\alpha_{n}}<c$
then with probability at least $1-n^{-1}$,
$\frac{1}{n}M_{\text{error}}\leq c^{-1}(2+\varepsilon)\frac{Kn_{\max}}{n_{\min}^{2}\lambda^{2}\alpha_{n}}$
Proof.
We now prove Theorem 3.3.
Under the model we have, we know that minimum eigenvalue of $B$ is $\lambda$. Use theorem B.5 to replace $h\left(\frac{2K}{c_{2}}\right)\left|\frac{1}{n}(M_{\text{error}})\right|$ with $h\left(\frac{2K}{c_{2}}\right)c^{-1}(2+\varepsilon)\frac{Kn_{\max}}{n_{\min}^{2}\lambda^{2}\alpha_{n}}$ in lemma 3.1.
We now have to show the existence of $c$ in Theorem B.5.
$\quad 2Kn_{\min}/n\geq c_{2}$
$\Rightarrow 1/n_{\min}^{2}\leq 4K^{2}/n^{2}c_{2}^{2}$
$\Rightarrow(2+\epsilon)\frac{Kn}{n_{\min}^{2}\lambda^{2}\alpha_{n}}\leq(2+\epsilon)\frac{4K^{3}}{n\lambda^{2}\alpha_{n}c_{2}^{2}}\leq(2+\epsilon)\frac{4K^{3}}{\lambda^{2}c_{2}^{2}}$
Let $c=(2+\epsilon)\frac{4K^{3}}{\lambda^{2}}c_{2}^{2}$
substitute $c$ to $h\left(\frac{2K}{c_{2}}\right)c^{-1}(2+\varepsilon)\frac{Kn_{\max}}{n_{\min}^{2}\lambda^{2}\alpha_{n}}$, we have that
|¯$|\bar{\mathcal{E}}(\mathcal{T})-\hat{\mathcal{E}}(\mathcal{T})|\leq h\left(\frac{2K}{c_{2}}\right)\frac{n_{\max}}{4c_{2}^{2}n_{\min}^{2}\alpha_{n}K^{2}}$
B.3.1 Proof of Corollary 3.3.1
Proof.
Now we prove Corollary 3.3.1. Note that $n\geq n_{\max}\geq n_{\min}\geq nc_{2}/2K$.
|¯$|\bar{\mathcal{E}}(\mathcal{T})-\hat{\mathcal{E}}(\mathcal{T})|\leq h\left(\frac{2K}{c_{2}}\right)\frac{n_{\max}}{4c_{2}^{2}n_{\min}^{2}\alpha_{n}K^{2}}\leq h\left(\frac{2K}{c_{2}}\right)\frac{1}{c_{2}^{4}\alpha n}$
Lemma B.6.
|^$|\mathcal{E}-\hat{\mathcal{E}}|\leq\sum_{j=1}^{K}\left(\left|\frac{p(C_{j})-\bar{p}(C_{j})}{p(C_{j})}\right|+\log\left(\frac{1}{p(C_{j})}\right)\left|p(C_{j})-\bar{p}(C_{j})\right|\right)+h\left(\frac{2K}{c_{2}}\right)\left|\frac{1}{n}(M_{\text{error}})\right|$
Proof.
First, we have that
|^||¯+¯^||¯|+|¯^||¯$|\mathcal{E}-\hat{\mathcal{E}}|\leq|\mathcal{E}-\bar{\mathcal{E}}+\bar{\mathcal{E}}-\hat{\mathcal{E}}|\leq|\mathcal{E}-\bar{\mathcal{E}}|+|\bar{\mathcal{E}}-\hat{\mathcal{E}}|\leq|\mathcal{E}-\bar{\mathcal{E}}|+h\left(\frac{2K}{c_{2}}\right)\left|\frac{1}{n}(M_{\text{error}})\right|$
Next,
$|\mathcal{E}-\bar{\mathcal{E}}|$ $\leq\left|\sum_{j=1}^{k}p(C_{j})\log(p(C_{j}))-\bar{p}(C_{j})\log(\bar{p}(C_{j}))\right|$
$\leq\left|\sum_{j=1}^{k}p(C_{j})\log(p(C_{j}))-\bar{p}(C_{j})\log(p(C_{j}))+\bar{p}(C_{j})\log(p(C_{j}))-\bar{p}(C_{j})\log(\bar{p}(C_{j}))\right|$
$\leq\sum_{j=1}^{k}\left|p(C_{j})\log(p(C_{j}))-\bar{p}(C_{j})\log(p(C_{j}))\right|+\left|\bar{p}(C_{j})\log(p(C_{j}))-\bar{p}(C_{j})\log(\bar{p}(C_{j}))\right|$
$\leq\sum_{j=1}^{k}\left|p(C_{j})-\bar{p}(C_{j})\right|\log\left(\frac{1}{p(C_{j})}\right)+\left|\frac{p(C_{j})-\bar{p}(C_{j})}{p(C_{j})}\right|$
Lemma B.7.
With probability at least $1-\frac{1}{n}$,
$\sum_{j=1}^{k}\left|p(C_{j})-\bar{p}(C_{j})\right|\leq K\sqrt{\frac{1}{2n}\log(2Kn)}$
Proof.
$\left|p(C_{j})-\bar{p}(C_{j})\right|=\frac{1}{n}\left|np(C_{j})-n_{j}\right|$
Now use Hoeffding bound, we notice that for any $j$
$\mathbb{P}(|n_{j}-np(C_{j})|\geq\delta)\leq 2\exp\left(-\frac{2\delta^{2}}{n}\right)$
Using union bound
$\mathbb{P}(\exists j\text{ such that }|n_{j}-np(C_{j})|\geq\delta)\leq\sum_{j=1}^{K}\mathbb{P}(|n_{j}-np(C_{j})|\geq\delta)\leq 2K\exp\left(-\frac{2\delta^{2}}{n}\right)$
$\exists j\text{ such that }|n_{j}-np(C_{j})|\geq\delta\Leftarrow\max|n_{j}-np(C_{j})|\geq\delta\Leftarrow\sum_{j=1}^{K}|n_{j}-np(C_{j})|\geq K\delta.$
Now, let $2K\exp\left(-\frac{2\delta^{2}}{n}\right)=\frac{1}{n}$, we have that $\delta=\sqrt{\frac{n}{2}\log(2Kn)}$
This gives us that with probability at least $1-\frac{1}{n}$,
$\sum_{j=1}^{k}\left|p(C_{j})-\bar{p}(C_{j})\right|\leq K\sqrt{\frac{1}{2n}\log(2Kn)}$
Lemma B.8.
With probability at least $1-\frac{1}{n}$
$n_{\min}\geq\frac{nc_{2}}{2K}$
where $c_{2}=2K\left(1-\sqrt{\frac{2\log(nK)}{np_{\min}}}\right)p_{\min}$ and $p_{\min}=\min\{p(C_{1})\dots p(C_{K})\}$
Proof.
Using the Chernoff inequality, we have
$\mathbb{P}\left(n_{j}\leq(1-\delta)np(C_{j})\right)\leq\exp\left(\frac{-np(C_{j})}{2}\right)$
Using the union bound
$\mathbb{P}(n_{\min}\leq nc_{2}/2K)\leq\mathbb{P}\left(\exists j\text{ such that }n_{j}\leq(1-\delta)np(C_{j})\right)\leq K\exp\left(\frac{-np_{\min}}{2}\right)$
Let $K\exp\left(\frac{-np_{\min}}{2}\right)=\frac{1}{n}$, we get $\delta=\sqrt{\frac{2\log(nK)}{np_{\min}}}.$ Finally, we have $c_{2}=2K\left(1-\sqrt{\frac{2\log(nK)}{np_{\min}}}\right)p_{\min}$
Appendix C Simulations
C.1 Hobby Examples
We can consider a list of things that a hypothetical individual "John" likes to do in his free time:
  • running / jogging
  • Drone flying / pilot Aerial drones
  • jazzercise / aerobics
  • making pottery / making ceramics
  • water gardening / aquatic gardening
  • caving / spelunking / potholing
  • cycling / bicycling / biking
  • reading
  • writing journals / journal writings/ journaling
  • sculling / rowing
C.2 Historical Examples
On the day December 3,
  • "Pope John X crowns Berengar I of Italy as Holy Roman Emperor"/ "Berengar I of Italy was crowned Emperor by Pope John X."
  • "USS Alfred becomes the first vessel to fly the Grand Union Flag; the flag is hoisted by John Paul Jones."/ "John Paul Jones hoisted the the Grand Union Flag on USS Alfred, the first vessel to fly Grand Union Flag."
  • "French General Jean Victor Marie Moreau decisively defeats the Archduke John of Austria near Munich in the Battle of Hohenlinden. "/"Archduke John of Austria was defeated by French General Jean Victor Marie Moreau near Munich in the Battle of Hohenlinden."
  • "Illinois becomes the 21st U.S. state."/ "Illinois joined U.S. as its 21st state."
  • "The Zollverein (German Customs Union) begins the first regular census in Germany"/"The first regular census was conducted by The Zollverein (German Customs Union) in Germany."
  • "The Duquesne Country and Athletic Club defeats an all-star collection of early football players 16–0."
  • "Following more than a month of Turkish–Armenian War, the Turkish-dictated Treaty of Alexandropol is concluded."/ "The Turkish-dictated Treaty of Alexandropol concluded after a month of Turkish–Armenian War."
  • "President Herbert Hoover delivers his first State of the Union message to Congress. It is presented in the form of a written message rather than a speech."/"President Herbert Hoover presents a written message as his first State of the Union message to Congress rather than a speech."
  • "The current flag of Singapore is adopted, six months after Singapore became self-governing within the British Empire."/"Singapore adopts it current flag six months after it self-govern within the British Empire."
  • "Ayatollah Ruhollah Khomeini becomes the first Supreme Leader of Iran.", "Iran has its first Supreme Leader: Ayatollah Ruhollah Khomeini."
Appendix D Prompt
This is the prompt we inserted for "Phi-3-mini-4k-instruct", "AI21-Jamba-1.5-Mini", "Cohere-command-r-08-2024".
    You are a expert in logical deduction and you are given 2 piece of texts: TEXT A and TEXT B.
    You are to identify if TEXT A implies TEXT B and TEXT B implies TEXT A at the same time.
    TEXT A:
    {text_A}
    TEXT B:
    {text_B}
    ## OUTPUT
    You are to return TRUE if TEXT A implies TEXT B and TEXT B implies TEXT A at the same time.
    otherwise, you are to return FALSE
This is the prompt we inserted for "Ministral-3B","Llama-3.3-70B-Instruct", "gpt-35-turbo"
    You are a expert in logical deduction and you are given 2 piece of texts: TEXT A and TEXT B.
    You are to identify if TEXT A implies TEXT B and TEXT B implies TEXT A at the same time.
    TEXT A:
    {text_A}
    TEXT B:
    {text_B}
    ## OUTPUT
    You are to return TRUE if TEXT A implies TEXT B and TEXT B implies TEXT A at the same time.
    otherwise, you are to return FALSE
    ##FORMAT:
    START with either TRUE or FALSE, then detail your reasoning
Appendix E Notes on Implementation in python
The default spectral clustering algorithm in python already implemented in k++ algorithm Choo et al. (2020).